Matt Gethers/CRI, Thailand/Labwork/Generating the HmgR Mutant/Week of 6.15.08

From OpenWetWare
Jump to: navigation, search

6.16.08

Today I need to take use Klenow to create blunt ends on both the NcoI-cut upstream fragment that I extracted from the gel on Friday and the SmaI-cut pUC18 that Sopapan gave me. I then need to ligate the two into each other. I have to check what protocol/enzymes they use to do ligations.

Before the Klenow work, I need to see if the PCR worked and repeat the digest. I have the PCR products from Friday afternoon (my second attempt - raised annealing temp to 55, added final extension time, included positive control), so I'll digest the 2724/1033 fragment like last Friday and run the product along with both PCR products out on a gel (1%, 45 minutes, 80 volts). Hopefully I'll have some bands to excise.

My PCR and Digest failed again. Time for some debugging. I'll ask Sopapan about her protocol to see where mine differs.

Reagents

It turns out that I'm using the wrong buffer - I should be using S (with Mg2+), not A. When I try the PCR again, I'll also make my own stock of the primers at 1 μM so I can use a larger volume.

Conditions

Sopapan does the following for the 1032/1033 PCR:

  • 95oC for 3 minutes


  • 95oC for 40 seconds
  • 55oC for 40 seconds
  • 72oC for 2 minutes/Kb

-

  • 72oC for 10 minutes

And she does the middle region for 35 cycles. Apparently PFU requires a bit longer time for denaturation and annealing.

I'm changing my protocol to reflect these insights and I've made notes on the latest run here.

If this amplification works, tomorrow morning I'll run the PCR again in a 50 μl reaction to produce enough to digest. I'll do the NcoI bit again and hopefully have something to excise and purify. At that point, I'll need to do the Klenow work, though I think I won't be using Klenow to do it :-)

6.17.08

To Do:

  • Run the two PCR products out on a gel to determine if the new protocol works.
    • If the PCR worked, set up a new 50 μl PCR using the 2724/1033 combination. (After I run the products out on the gel, if 10 μl if easy to visualize, I won't have to make more for the digest).
    • If both PCRs failed, check once more with Sopapan if I'm doing everything as she did, then talk with Mayuree.
    • If the positive control worked but the 1033/2724 failed again, maybe gradient PCR for annealing temp?
  • Assuming the Upstream framgent PCR works, digest the product with NcoI, run out on a gel and excise the correct band.
  • Do a gel clean up of the excised bands.

That should keep me busy for one day. I'll worry about the blunting and ligation tomorrow if all of the above goes well.

I'm running a gel with both the 1032/1033 and 1033/2724 PCR products from yesterday afternoon (using buffer S and other modifications).

Now I'm running a gel with the 1033/2724 and 2736/2737 PCR products and the 1033/2724, 2736/2737, and pET-11a digests (the latter two are for the HmgA purification project).

6.18.08

To Do:

  • Ask Mayuree how to select for polarity of insert.
  • Use End-It kit to fill in upstream fragment and pUC18-SmaI vector.
  • Ligate fragments 1 and 3 into pUC18-SmaI and pUC18-PstI respectively.


Summary:

Did End-IT kit on both the downstream HmgA fragment (1 by Sopapan's numbering) and pUC18-SmaI. Due to the pleasant surprise of the HmgR ORF amplification working, I didn't have time to do the ligation, so that will be the first thing I do tomorrow.

6.19.08

  • Ligate fragments 1 and 3 into pUC18-SmaI and pUC18-PstI respectively.
  • Transform into some cloning strain?
  • How to select for polarity of insert?

Summary:

I forgot that I had the cultural orientation today, so everything is postponed until Friday.

6.20.08

  • Ligate fragments 1 and 3 into pUC18-SmaI and pUC18-PstI respectively.
  • Transform into DH5 α.
  • How to select for polarity of insert?
    • Will talk with Mayuree about this later, but pUC18 does allow for Blue-White selection, so that's a plus.

Summary:

Ligation done. Transformation into DH5 α done as well. All plates in 37oC incubator at 16:45.